# When Speculation Spills Secrets: Side Channels via Speculative Decoding in LLMs

### v3

@@ 4.5 Results
Table 2 shows the results for the Query Fingerprinting Attack on LADE, REST, and BiLD.
Exact Knowledge Scenarios. In both Experiment 1 and 2, the attacker leaks the input prompt with high accuracy: with accuracy of up to 70.8%, 100%, and 77.6% for LADE, REST, and BiLD respectively, compared to random guessing (2% for a set of 50 prompts). In Experiment 1, where we study randomly selected prompts and the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD. In Experiment 2, where all our prompts have a similar structure but different semantic meaning, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 70.8%, 99.6%, and 77.6% respectively. This indicates that our attack success rate is not influenced by the structure of the prompt, but relies and leaks the semantic meanings of the prompt and response.
@@ 4.6 Ablation Studies
Table  2 presents an ablation study for the attack accuracy as the temperature of our target model varies. For LADE and BiLD, as the temperature increases from 0.3 to 1.0, the attack accuracy for both Experiment 1 and 2 decreases. This is because at lower temperatures, the generations and speculation patterns are more stable and alike. Hence, the output tokens for the same prompt and the packet sizes are more similar across runs, making the attack more successful. For REST, the attack accuracy is stable for both Experiment 1 and 2 despite the increasing temperature. This is because the features are so pronounced that it withstands perturbations in the output caused by the increasing temperature.
@@ 4.7 Attack on Remote vLLM Server
To show real-world practicality, we demonstrate our attack on a remote LLM inference server on vLLM, with Llama3 8B Instruct as the target model and EAGLE speculative decoding. Our LLM server is hosted on a cloud-based A100 GPU, while our client is located in an academic network more than 1500 miles away. Our attacker snoops the encrypted network packets in between the server and the client, using Wireshark, and captures the bytes per packet. We also compare our attack with prior work, Carlini & Nasr (2024), that used packet inter-arrival-times as a side-channel, that varies based on inference-time optimizations. To mimic their setup, which focused on ChatGPT, that sends tokens generated in the same iteration in separate packets, we modify vLLM to similarly send separate packets for individual tokens generated in the same iteration. For our packet size side channel, we generate our trace by combining sizes of packets received within 55 ms of each other (based on observed network latencies).
Table 3 shows that, under no additional server load, our attack achieves a maximum of 77.6% accuracy, significantly outperforming prior work which achieves a maximum accuracy of 14.4%. This is because inter-arrival times provide a noisy signal due to varying network contentions. Moreover, under spikes in server load, the inter-arrival times (time per output token) can significantly spike as observed in recent work Zhang et al. (2024b), causing test-time observations to vary dramatically compared to training time. Consequently, assuming a standard log-normal distribution of time-per-iteration to mimic high server load, the accuracy of prior timing channel attacks drops to 6.4%. In comparison, our attack accuracy is unaffected, since we rely on the packet size side-channel, which remains robust to variations in server loads or network congestion.
@@
Temperature $\rightarrow$ 0.3 0.6 0.8 1.0 0.3 0.6 0.8 1.0
Packet Size Side Channel (ours) 77.6% 65.6% 44.8% 24.0% 77.6% 65.6% 44.8% 24.0%
Inter-Arrival Time Side Channel 14.4% 6.0% 5.2% 4.8% 6.4% 4% 2.4% 2.8%
(by Carlini & Nasr (2024))
Table 3: Accuracy of query fingerprinting attack (experiment 1) on vLLM with EAGLE’s speculative decoding at different temperatures (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training. We evaluate the attack using our packet-size side channel, and with inter-arrival times based side channel from prior work Carlini & Nasr (2024).
@@ 4.8 Out-of-Distribution Training
To study the effect of distribution mismatch between profiling and attack phases, we conducted an additional experiment where the attacker trains on an out-of-distribution (OOD) set of queries. Shown in  A.5, we used GPT-4o to generate 50 common diseases that users typically ask AI chatbots about, prepending the phrase “what are the symptoms of” to each. This OOD set was used for training, while evaluation was performed on the same 50 disease-related queries from Experiment 1 (with model temperature of 0.3 to 0.8), ensuring no prior knowledge of the test dataset during training.
Despite the lack of prior knowledge of the true query set, the attack substantially outperformed random guessing ($\sim$2%). LADE achieved 23% - 25% accuracy, REST 35% - 36%, and BiLD 23% - 25%, where success was defined as predicting a disease with similar symptoms (as per GPT-5 mini) compared to the ground truth, as temperature of the target model varies between 0.3 and 0.8 (see Table  7 in Appendix E for detailed results). These results show that although accuracy drops relative to the in-distribution setting, speculative decoding still leaks strong and learnable signals even under OOD conditions, without being much affected by temperature, underscoring the robustness of our side-channel attack.
@@ 5 Leaking Private Data Used for Speculation
Speculative decoding techniques may use carefully tuned data to ensure high correct speculation rates, which may contain private user data or other intellectual property. In this section, we show how a malicious user can leak out private data used for predictions. We show this on REST He et al. (2024), which relies on retrieval from a datastore to generate speculative tokens. This datastore can be populated with proprietary information to tailor model output to specific domains, or data collected from user interactions, which may be leaked by our attack. A similar attack can also leak speculation hyperparameters, which can be a company’s intellectual property, as shown in Appendix C.
@@
2. Leveraging Common Words. The LLM is prompted to generate paragraphs containing the most frequently occurring words, drawn from a dataset of the 10,000 most common English words Brants & Franz (2021). This exploits the high likelihood that phrases in the datastore contain these words.
@@ 6.1 Mitigating Query Fingerprinting Attacks
No Padding Fixed Padding Random Padding (bytes) Packet Size Overhead
(1024 bytes) D $\rightarrow$ 6 12 24 48 D $\rightarrow$ 6 12 24 48
Experiment 1
@@
[…]
Experiment 2
[…]
Experiment 3
[…]
Table 4: Attack accuracy under constant size padding (packets padded to 1024 bytes) and variable-length padding (random padding) mitigations. Variable-length padding follows a uniform distribution, $\epsilon\sim Unif(0,D)$, with $D$ from 6 to 48. (Attack: 20 traces per prompt; model temperature: 0.8).
[…]
2. Reduce Granularity of Attacker Observation: Alternatively, the LLM server can aggregate output tokens over multiple iterations and return them together to the client. This limits observability of fine-grain speculation patterns from packet sizes, reducing the success of fingerprinting attacks.
Table 5 shows the impact of aggregating output tokens over 3 to 20 iterations. The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drops to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating tokens over 20 iterations down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20%. Thus, token aggregation is an effective mitigation. A downside of aggregating tokens over larger iterations is the increased latency between outputs, which can hinder system responsiveness. At the extreme, aggregating all tokens in a single packet resembles a non-streaming LLM.
Experiment 1 Experiment 2 Experiment 3
AGG $\rightarrow$ 1 3 5 10 20 1 3 5 10 20 1 3 5 10 20
LADE 65.6% 53.2% 40% 24.4% 10.8% 64.8% 45.2% 28.8% 16.4% 9.2% 13.6% 13.2% 14.4% 9.2% 5.2%
REST 100% 96.8% 89.2% 80.4% 52% 100% 97.6% 92.8% 75.6% 52% 40% 33.2% 28.8% 20.4% 17.6%
BiLD 49.2% 34% 30% 19.2% 11.2% 75.2% 57.6% 41.6% 30% 16.8% 20.4% 15.6% 10% 10% 4%
Table 5: Token Aggregation Mitigation. Attack accuracy (20 traces/prompt, temperature of 0.8) as aggregation granularity (AGG), i.e., iterations tokens are aggregated over, varies from 3 to 20.
@@ C.3 Leaking Hyper-parameter G in LADE
Appendix D Impact of Mitigation for Token-Length Side Channel
In this experiment, we want to measure the impact on our attacks of mitigation against prior attacks, such as token-length side channels Weiss et al. (2024) that leaks the value of a token based on its associated character length. We evaluate the impact of a mitigation, where every token has been padded to the same size to mask its character length. We measure the attack success rate for our attacks with this mitigation. Note that even with this mitigation, there is variation in the packet sizes, due to simply the number of tokens contained in each packet, which varies due to speculation. Thus, measuring the packet sizes in this scenario is the same as measuring the token counts.
@@ Appendix D Impact of Mitigation for Token-Length Side Channel
Temperature
0.3 0.6 0.8 1.0
Experiment 1
LADE 51.6% 38% 27.6% 22.8%
REST 24% 17.2% 10.8% 8.4%
BiLD 67.6% 43.2% 30.4% 22.8%
Experiment 2
LADE 41.2% 24.8% 22.4% 11.2%
REST 25.6% 14% 14.4% 10%
BiLD 82.4% 62.4% 46.8% 33.2%
Experiment 3
LADE 8% 6.8% 6.8% 3.2%
REST 6.4% 4.8% 8% 5.2%
BiLD 8.4% 10% 10.8% 6.8%
Table 6: Accuracy of the Query Fingerprinting Attack with each token padded to the maximum token size (1024 bytes)
Appendix E Out of Distribution Training

### v4

@@ 4.5 Results
Figure 3 shows the results for the query fingerprinting attack on LADE, REST, and BiLD.
Exact Knowledge Scenarios. In both Experiment 1 and 2, the attacker leaks the input prompt with high accuracy: with accuracy of up to 70.8%, 100%, and 77.6% for LADE, REST, and BiLD respectively, compared to random guessing (2% for a set of 50 prompts). In Experiment 1, where the set of profiled and leaked prompts are identical, the attack success rates improve as the number of traces per query (TPQ) profiled increases, reaching up to 68% accuracy and an F1 score of 0.66 with 30 TPQ for LADE, reaching 99.6% accuracy with 30 TPQ and an F1 score of 1.0 for REST, and reaching 52% accuracy with 30 TPQ and an F1 score of 0.55 for BiLD. In Experiment 2, where all our prompts have a similar structure but different semantic meaning, the results are similar, with attack accuracy for LADE, REST, and BiLD reaching 90.8%, 99.6%, and 95.2% respectively. This indicates that our attack success rate is not influenced by the structure of the prompt, but relies and leaks the semantic meanings of the prompt and response.
@@ 4.6 Ablation Studies
Figure 4 presents an ablation study for the attack accuracy as the temperature of our target model varies. For LADE and BiLD, as the temperature increases from 0.3 to 1.0, the attack accuracy for both Experiment 1 and 2 decreases. This is because at lower temperatures, the generations and speculation patterns are more stable and alike. Hence, the output tokens for the same prompt and the packet sizes are more similar across runs, making the attack more successful. For REST, the attack accuracy is stable for both Experiment 1 and 2 despite the increasing temperature. This is because the features are so pronounced that it withstands perturbations in the output caused by the increasing temperature.
Figure 4: Attack accuracy of query fingerprinting as temperatures vary (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
@@ 4.7 Attack on Remote vLLM Server
To show real-world practicality, we demonstrate our attack on a remote LLM inference server on vLLM, with Llama3 8B Instruct as the target model and EAGLE speculative decoding. Our LLM server is hosted on a cloud-based A100 GPU, while our client is located in an academic network more than 1500 miles away. Our attacker snoops the encrypted network packets in between the server and the client, using Wireshark, and captures the bytes per packet. We also compare our attack with prior work, (Carlini & Nasr, 2024), that used packet inter-arrival-times as a side-channel, that varies based on inference-time optimizations. To mimic their setup, which focused on ChatGPT, that sends tokens generated in the same iteration in separate packets, we modify vLLM to similarly send separate packets for individual tokens generated in the same iteration. For our packet size side channel, we generate our trace by combining sizes of packets received within 55 ms of each other (based on observed network latencies).
@@
[…]
Random Forest Packet Size (ours) 77.6% 65.6% 44.8% 24.0% 77.6% 65.6% 44.8% 24.0%
Inter-Arrival Time 14.4% 6.0% 5.2% 4.8% 6.4% 4% 2.4% 2.8%
(Carlini & Nasr, 2024)
GMM Packet Size (ours) 27.2% 7.2% 4.8% 4% 27.2% 7.2% 4.8% 4%
Inter-Arrival Time 2.6% 1.2% 3.2% 1.2% 1.6% 1.2% 2.8% 1.2%
(Carlini & Nasr, 2024)
Table 1: Attack accuracy for the query fingerprinting (Experiment 1 - Exact Knowledge) on vLLM with EAGLE’s speculative decoding at different temperatures (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training. We evaluate the attack using our packet-size side channel, and with inter-arrival times based side channel from prior work (Carlini & Nasr, 2024).
Table 1 shows that, under no additional server load, our attack achieves a maximum of 77.6% accuracy, significantly outperforming prior work which achieves a maximum accuracy of 14.4% (both using Random Forest classifier). This is because inter-arrival times provide a noisy signal due to varying network contentions. Moreover, under spikes in server load, the inter-arrival times (time per output token) can significantly spike as observed in recent work (Zhang et al., 2024b), causing test-time observations to vary dramatically compared to training time. Consequently, assuming a standard log-normal distribution of time-per-iteration to mimic high server load  (Benson et al., 2009), the accuracy of prior timing channel attacks drops to 6.4%. In comparison, our attack accuracy is unaffected, since we rely on the packet size side-channel, which remains robust to variations in server loads or network congestion. Even when using GMM instead of Random Forest like prior work (Carlini & Nasr, 2024), the trends are similar, with our packet-size side-channel outperforming prior inter-arrival time side-channel (Carlini & Nasr, 2024) under both low and high server loads.
@@ 4.8 Out-of-Distribution Training
To study the effect of distribution mismatch between profiling and attack phases, we conducted an additional experiment where the attacker trains on an out-of-distribution (OOD) set of queries. Shown in Appendix A.5, we used GPT-4o to generate 50 common diseases that users typically ask AI chatbots about, prepending the phrase “what are the symptoms of” to each. This OOD set was used for training, while evaluation was performed on the same 50 disease-related queries from Experiment 1 (with model temperature of 0.3 to 0.8), ensuring no prior knowledge of the test dataset during training.
Despite the lack of prior knowledge of the true query set, the attack substantially outperformed our previous naive random guessing attacks ($\sim$2%). LADE achieved 23% - 25% accuracy, REST 35% - 36%, and BiLD 23% - 25%, where success was defined as predicting a disease with similar symptoms (as per GPT-5 mini) compared to the ground truth, as temperature of the target model varies between 0.3 and 0.8 (see Table  3 in Appendix  F for detailed results). For comparison, we also perform a more “sophisticated” random guessing attack using OOD data: for each query in our test set, we randomly pick a disease from the OOD set as our guess, and check whether the guess has symptoms overlapping with the ground truth. This random guessing with OOD data has 6% accuracy (higher than 2% with our previous naive random guessing); this is still much lower than the 23% to 35% accuracy we achieve by training our classifier using the speculation traces observed with OOD data.
These results show that although accuracy drops relative to the in-distribution setting, speculative decoding still leaks strong and learnable signals even under OOD conditions, without being much affected by temperature, underscoring the robustness of our side-channel attack.
@@ 5 Leaking Private Data Used for Speculation
Speculative decoding techniques may use carefully tuned data to ensure high correct speculation rates, which may contain private user data or other intellectual property. In this section, we show how a malicious user can leak out private data used for predictions. We show this on REST (He et al., 2024), which relies on retrieval from a datastore to generate speculative tokens. This datastore can be populated with proprietary information to tailor model output to specific domains, or data collected from user interactions, which may be leaked by our attack. A similar attack can also leak speculation hyperparameters, which can be a company’s intellectual property, as shown in Appendix C.
@@
2. Leveraging Common Words. The LLM is prompted to generate paragraphs containing the most frequently occurring words, drawn from a dataset of the 10,000 most common English words (Brants & Franz, 2021). This exploits the high likelihood that phrases in the datastore contain these words.
@@ 6.1 Mitigating Query Fingerprinting Attacks
B. Random Padding. Packets can also be padded with a random number of bytes ($\epsilon$) selected from a uniform distribution of $\epsilon\sim Unif(0,D)$, where D is an upper bound of padding and $\epsilon$ is an integer.
@@
Figure 6: Attack accuracy under variable-length padding (random padding) added to payloads within packets as a mitigation. Variable-length padding follows a uniform distribution, $\epsilon\sim Unif(0,D)$, with $D$ from 6 to 48. (Attack: 20 traces per prompt; model temperature: 0.8).
2. Reduce Granularity of Attacker Observation: Alternatively, the LLM server can aggregate output tokens over multiple iterations and return them together to the client. This limits observability of fine-grain speculation patterns from packet sizes, reducing the success of fingerprinting attacks. However, token aggregation linearly increases the inter-arrival time for packets, which linearly increases the latency observed by a user between text renderings that may hinder user experience.
Figure 7 shows the accuracy for all our fingerprinting attacks after aggregating output tokens over 3 to 20 iterations. The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drop to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating tokens over 20 iterations, down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20% respectively. Thus, token aggregation is an effective mitigation. A downside of aggregating tokens over larger iterations is the increased latency between outputs, which can hinder system responsiveness. At the extreme, aggregating all tokens in a single packet resembles a non-streaming LLM.
Figure 7: Token Aggregation Mitigation. Attack accuracy (30 traces/prompt, temperature of 0.8) decreases as aggregation granularity (AGG) increases, i.e., iterations tokens are aggregated over, varies from 1 (no aggregation) to 20.
3. Packet Splitting. Another way of limiting attacker observability of token counts, is to split the multiple tokens generated in an iteration into individual packets that are spread out in time throughout the decode iteration time period. To ensure that the attacker cannot observe even an approximate count of tokens-per-iteration by tracking the number of packets sent during one decode iteration, the defense needs to further also limit the number of packets per iteration and carry excess tokens to later iterations. This can introduce an overall latency impact, varying based on the framework, model and speculation technique. Future works can develop novel speculation techniques with such traffic shaping built in, to get both security and performance.
@@ C.3 Leaking Hyper-parameter G in LADE
Appendix D Results for Random Padding Mitigation
@@
No Padding Fixed Padding Random Padding (bytes) Payload Size Overhead
(1024 bytes) D $\rightarrow$ 6 12 24 48 D $\rightarrow$ 6 12 24 48
Experiment 1 - Exact Knowledge
[…]
Experiment 2 - Exact Knowledge (Constrained)
[…]
Experiment 3 - Approx Knowledge
[…]
Table 2: Attack accuracy under constant size padding (payload within packets padded to 1024 bytes) and variable-length padding (random padding) mitigations. Variable-length padding follows a uniform distribution, $\epsilon\sim Unif(0,D)$, with $D$ from 6 to 48. (Attack: 20 traces per prompt; model temperature: 0.8).
Appendix E Impact of Mitigation for Token-Length Side Channel
In this experiment, we want to measure the impact on our attacks of mitigation against prior attacks, such as token-length side channels (Weiss et al., 2024) that leaks the value of a token based on its associated character length. We evaluate the impact of a mitigation, where every token has been padded to the same size to mask its character length. We measure the attack success rate for our attacks with this mitigation. Note that even with this mitigation, there is variation in the packet sizes, due to the number of tokens contained in each packet, which varies due to speculation. Thus, measuring the packet sizes in this scenario is the same as measuring the token counts.
Figure 10 shows the accuracy for our attack despite this mitigation as temperature varies. At smaller temperatures, the accuracy of our attacks is still significant (with temperature 0.3, in both exact knowledge attacks (experiment 1 and 2), for LADE, the accuracy is about 40-50%, and for BiLD, it is about 80%). At high temperatures, our attack accuracy is reduced: with temperature 1.0, in both the exact knowledge scenarios (experiment 1 and 2), for LADE, the accuracy is reduced to around 20%, and for BiLD, it is reduced to 30%.
Figure 10: Accuracy of the query fingerprinting attack with each token padded to the maximum size of tokens that can be generated per iteration (1024 bytes)
[…]
Appendix F Out of Distribution Training
F.1 Attack Accuracy
Table 3 shows the attack accuracy using out of distribution dataset with 50 queries about diseases as training data, and using our original evaluation data set from Experiment 1 as the test set. Our attack achieves an accuracy of 23% to 36% of guessing overlapping symptoms by training wiht the OOD set, achieving significantly higher accuracy than random guessing using OOD set, which achieves only a accuracy of 6%.
@@ F.1 Attack Accuracy
F.2 Degree of Overlap Between OOD Training Set and Evaluation Set
Since we use a 50 class OOD data set (generated using ChatGPT) as training set, and a 50 class data set for the evaluation set (used in Experiment 1), there is no direct 1-1 correspondence between training and evaluation sets. However, artificially, if the symptoms of each disease in the OOD training set has a high overlap with all the diseases of the test set, then our attack accuracy can be artificially high. To understand whether this is the case, we study the the overlap in symptoms between diseases in the OOD training set and that in the evaluation set (Experiment 1 - Exact Knowledge).
Figure 13 shows a 2D matrix illustrating symptom overlap between diseases in the OOD training set (X-axis) and diseases in the evaluation set from Experiment 1 (Y-axis). Both sets have 50 diseases. A cell is colored white if the corresponding pair of diseases shares overlapping symptoms. To determine overlap, we query an LLM (e.g., ChatGPT) with the two diseases and check whether they have common symptoms.
In the OOD training set, a small number of diseases (3-4) share symptoms with multiple (5-6) diseases in the evaluation set, and vice versa. However, most training-set diseases overlap with only one or two evaluation-set diseases. Additionally, some evaluation-set diseases have no symptom overlap with any training-set disease; in such cases, an attacker cannot infer the patient’s symptoms. This is expected, as the OOD training set is independently constructed from the top 50 diseases most commonly queried on ChatGPT, while the evaluation set is derived from real-world user queries to medical chatbots.
This diversity in overlap patterns reflects a realistic deployment setting. Consequently, the observed attack accuracy of 23%–36% is realistic.
In comparison, a random guessing baseline attack, that assigns a random OOD training-set disease to each evaluation set disease achieves only 6% accuracy, aligning with our observation that most diseases in OOD training set do not overlap in symptoms with many test set diseases. This highlights the potency of our attack using speculative decoding when realistic OOD data is used for training.
Figure 11: Overlap in symptoms between diseases in the OOD Training set (X-axis) and in the Evaluation set used in Experiment 1 - Exact Knowledge (Y-axis). A white entry indicates that two diseases have similar symptoms (as per GPT-4o), while a black entry indicates no overlap. Overall, there is minimal overlap in symptoms across any random pair of diseases from OOD training dataset and evaluation dataset, representative of realistic OOD datasets.
Appendix G Comparing Classification Methods (Random Forest, GMM, CNN)
In this section, we evaluate the accuracy of the query fingerprinting attack (Experiment 1 - Exact Knowledge) using different classifiers - Random Forest, GMM and CNN. We study this with for the three speculative decoding techniques LADE, REST and BiLD for a range of temperatures (0.3, 0.6, 0.8, 1), all with traces per query (TPQ) of 30.
For the GMM classifier, we trained the scikit-learn’s GMM with 1 component per class, and make predictions via Bayes’ Rule. For the CNN classifier, we use three 1D convolution layers with max pooling followed by average pooling and a fully connected layer. For the Random Forest classifier, our setup is similar to Section 4.4.
Figure 12: Accuracy of query fingerprinting attack (Experiment 1 - Exact Knowledge) using Random Forest, GMM, and CNN classifiers, for temperatures 0.3, 0.6, 0.8, 1.0 (using 30 traces per query).
As shown in Figure 12, the attack accuracy for the GMM classifier ranges from 13-54% for LADE, 79-96% for REST, and 14-51% for BiLD. However the GMM performs worse than the Random Forest classifier which achieves an accuracy of 61-92% for LADE, 99-100% for REST, and 34-74% for BiLD. While prior work Carlini & Nasr (2024) showed that GMM can achieve query fingerprinting with high accuracy (up to 100%), we observe that their attacks were restricted to 2 classes, contrary to our attacks that classify across 50 classes, in which scenario the GMM achieves lower accuracy and the Random Forest is able to outperform it.
In comparison, the CNN performs worse than both the Random Forest and the GMM, achieving an accuracy of 10-22% for LADE, 52-80% for REST, and 5-12% for BiLD. This is because the training data set is relatively small, with only 250 to 1500 training data points. So we observe that the CNN overfits to a given trace, leading to lower accuracy.
These results lead us to choose Random Forest classifier for all our attacks in this paper, given that it outperforms GMM and CNN.
Appendix H Scaling Behavior
In this section, we study how the accuracy of the query fingerprinting attack (with Experiment 1 - Exact Knowledge) with temperature of 0.8 changes as the dataset size (number of diseases) scales from 5 to 50. As shown in Table 4, for BiLD and REST, the accuracy remains relatively stable as the number of diseases increase. The attack accuracy for LADE decreases by about 7 percent when we double the dataset size from 25 diseases to 50. This shows that the attack is relatively robust to scaling of the dataset size.
Number of Diseases in Dataset
5 10 15 20 25 30 35 40 45 50
LADE 92% 86% 89.3% 84% 76.8% 74.7% 74.3% 73.5% 74.2% 68%
REST 100% 100% 100% 100% 100% 100% 100% 99.5% 99.6% 99.6%
BiLD 48% 40% 60% 58% 45.6% 48% 49.1% 48.5% 48.8% 52%
Table 4: Accuracy of query fingerprinting attack (Experiment 1 - Exact Knowledge) with temperature of 0.8, using 30 traces per query (TPQ) for training as the number of diseases in the dataset scales from 5 to 50.
Next, we study how the training cost scales with the dataset size. Specifically, we measure the cost in terms of the number of traces per query required to reach 90% of the fingerprinting attack (Experiment 1 - Exact Knowledge, with temperature 0.8, and dataset of 50 diseases). We measure this cost as the dataset size scales from 5 to 50 diseases.
As shown in Table 5, we observe that the number of traces (cost) remains relatively stable (for REST) or increases moderately (LADE and BiLD). For REST, this is because the retrieval based speculation provides an extremely stable signal, which allows for high accuracy in the attack with a limited number of traces. In comparison, both LADE and BiLD have more variations, requiring more traces per query as the number of diseases increase.
Number of Diseases in Dataset
5 10 15 20 25 30 35 40 45 50
LADE(61.2%) 3 3 3 4 5 7 7 10 9 9
REST(89.6%) 4 2 3 4 4 5 5 3 4 4
BiLD(46.8%) 22 30 15 16 28 29 28 27 19 19
Table 5: Number of queries needed to achieve 90% of the fingerprinting attack accuracy (exact knowledge - experiment 1) with temperature of 0.8, as the number of diseases in the dataset scales from 5 to 50.
Appendix I Ablation: Query Fingerprinting Attack on Insurance Chatbot
To show that our attack is not restricted to the medical domain, and also works on other domains, we demonstrate our attack on an insurance chatbot.
Figure 13: Attack accuracy of query fingerprinting with 50 insurance-related queries as temperatures vary (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.
We perform the query fingerprinting attack over a set of 50 queries consisting of insurance-related questions, generated by asking ChatGPT for commonly asked insurance-related questions by users. These include questions on health, car, and home insurance. Learning whether a user is asking such questions can allow an attacker to know private information such as whether a user owns a car or a home or has health concerns. We perform the Exact Knowledge, Exact (Constrained) and Approx Knowledge experiments , as designed for the medical chatbot in 4.3. Our REST and BiLD, our attack on an insurance chatbot achieves 97.2–100% and 37.6–91.6% query-fingerprinting accuracy for temperatures 0.3–1.0, slightly outperforming the results we observed on the medical chatbot. LADE with 17.2–59.2% attack accuracy performs worse than in the medical case, however it still outperforms random chance (2%). This shows that the attack does not rely on any medical specific terminologies and can generalize to more diverse domains.